# GMLM: Bridging Graph Neural Networks and Language Models for Heterophilic Node Classification

### v4

@@ 4.4 Bi-Directional Attention Fusion and Classification
$\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$ (9)
@@
$\mathbf{h}_{concat,i}=[\mathbf{h}_{G\to T,i}\,;\,\mathbf{h}_{T\to G,i}]\in\mathbb{R}^{2\cdot d_{PLM}}$ (10)
@@
$\mathbf{h}_{fused,i}=\text{FusionNet}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$ (11)
@@
$\hat{\mathbf{y}}_{i}=\text{Softmax}(\text{MLP}_{classifier}(\mathbf{h}_{fused,i}))$ (12)
@@
5 Training Procedure
The model training proceeds in two distinct stages: (1) GNN Contrastive Pre-training, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. This two-stage approach allows the GNN to develop a strong structural foundation before being integrated with the powerful but computationally intensive language model. We also provide an analysis of the impact on embeddings across different stages of training in LABEL:app:imp-train.
5.1 Stage 1: GNN Contrastive Pre-training
@@
5.1.1 Data Augmentation via Soft Masking
@@ 5.1.2 Contrastive Objective
$\mathcal{L}_{pretrain}=\mathcal{L}_{NT-Xent}(\mathbf{H}_{G,1},\mathbf{H}_{G,2},\tau_{pretrain})$ (13)
@@
5.2 Stage 2: End-to-End Fine-tuning for Node Classification
@@
5.2.1 Active Node Selection and GNN Perturbation
@@
5.2.2 Supervised Classification Loss
@@
5.2.3 Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated through the entire model. All trainable parameters are updated using the AdamW optimizer Loshchilov and Hutter (2019) with differential learning rates:
@@ 5.2.3 Optimization and Hyperparameters
H2GCNZhu et al. (2020b) 69.76$\pm$3.0 79.09$\pm$3.5 80.18$\pm$1.9 70.73$\pm$0.9 47.09$\pm$0.3
FAGCNBo et al. (2021) 76.43$\pm$3.1 84.55$\pm$4.8 83.16$\pm$1.4 75.58$\pm$0.5 49.83$\pm$0.6
JacobiConvWang and Zhang (2022) 73.57$\pm$4.3 81.80$\pm$4.1 76.31$\pm$11.3 73.81$\pm$0.3 49.43$\pm$0.5
GBK-GNNDu et al. (2022) 66.19$\pm$2.8 80.00$\pm$3.0 72.98$\pm$3.3 72.49$\pm$1.0 44.90$\pm$0.3
OGNNWang et al. (2022) 71.91$\pm$1.8 85.00$\pm$2.3 79.30$\pm$2.1 72.08$\pm$2.4 47.79$\pm$1.6
SEGSL 66.67$\pm$4.1 85.00$\pm$2.0 79.30$\pm$1.8 72.73$\pm$0.8 47.38$\pm$0.2
DisamGCLZhao et al. (2024) 50.48$\pm$2.0 65.00$\pm$1.2 57.89$\pm$0.0 67.78$\pm$0.3 43.90$\pm$0.4
LLM4HeG Wu et al. (2025) (fine-tuned LLM/SLMs and distilled SLMs)
Vicuna 7B 77.62$\pm$2.9 89.09$\pm$3.3 86.14$\pm$2.1 76.82$\pm$0.5 51.53$\pm$0.4
Bloom 560M 75.48$\pm$2.1 80.00$\pm$4.0 86.49$\pm$1.9 76.16$\pm$0.6 51.52$\pm$0.5
Bloom 1B 75.71$\pm$1.4 83.86$\pm$2.8 83.86$\pm$1.7 74.99$\pm$0.5 52.33$\pm$0.6
7B-to-560M 75.00$\pm$4.0 88.18$\pm$2.2 87.19$\pm$2.5 75.78$\pm$0.2 51.51$\pm$0.4
7B-to-1B 77.38$\pm$2.7 88.18$\pm$4.0 86.14$\pm$1.5 75.37$\pm$0.9 51.58$\pm$0.4
@@
GTE-Base 74.25$\pm$2.5 85.90$\pm$4.9 89.81$\pm$2.5 73.89$\pm$2.1 50.64$\pm$0.7
Snowflake-embed 79.16$\pm$4.8 97.18$\pm$2.9 92.04$\pm$1.6 76.81$\pm$0.9 54.63$\pm$0.4
Table 1: Accuracy for node classification of different methods. Bold denotes the best score while underline denotes the second best.
6 Results
We evaluate GMLM on five widely-used heterophilic benchmark datasets. Following standard protocols, we randomly split the nodes into train, validation, and test sets with a proportion of 48%/32%/20% for the Cornell, Texas, Wisconsin, and Actor datasets, and 50%/25%/25% for Amazon following Platonov et al. (2024). All reported results in Table 1 are the mean accuracy (%) and standard deviation averaged over 10 independent runs.
@@ 6 Results
A comparison of our model variants reveals that performance scales with the representational power of the PLM, as Snowflake-embed (305M) consistently outperforms GTE-Base (110M). However, even the smaller GMLM-GTE-Base is highly competitive, outperforming the majority of all baselines and securing the second-best result on Wisconsin. This confirms the robustness and effectiveness of the core GMLM framework itself.
7 Conclusion
@@
In this paper, we introduced the Graph Masked Language Model (GMLM), a novel framework that effectively and efficiently integrates deep graph structural learning with the rich semantic understanding of Pre-trained Language Models. By employing a two-stage training procedure featuring a contrastive GNN pre-training phase and an end-to-end fine-tuning stage with dynamic active node selection, our model addresses key challenges of scalability and robust representation learning. The use of a sophisticated bi-directional cross-attention module for fusion proved critical, enabling GMLM to achieve new state-of-the-art results on four out of five widely-used heterophilic benchmarks. Our model significantly outperforms both specialized GNNs and much larger LLM-based approaches, underscoring that a carefully designed, deeply integrated architecture can be more powerful and capital-efficient than brute-force scaling for text-rich graph representation learning.
8 Limitations and Future Work
@@
Appendix A Dataset Statistics
Table 2 provides an overview of the key statistics for the datasets used in our experiments from HeTGB Li et al. (2025).
Name Nodes Edges H($G$) Classes
Cornell 195 304 0.13 5
Texas 187 328 012 5
Wisconsin 265 530 0.20 5
Actor 4416 12172 0.56 5
Wisconsin 24492 93050 0.38 5
Table 2: Statistics of the heterophilic benchmark datasets used.
Appendix B LLM Usage
We used the Gemini2.5 Pro models to help refine our grammar and paper presentation.

### v6

@@ 4.4. Bi-Directional Attention Fusion and Classification
(8) $\text{Attention}(\mathbf{Q},\mathbf{K},\mathbf{V})=\text{softmax}\left(\frac{\mathbf{Q}\mathbf{K}^{T}}{\sqrt{d_{k}}}\right)\mathbf{V}$
@@
(9) $\mathbf{h}_{concat,i}=[\mathbf{h}_{G\to T,i}\,;\,\mathbf{h}_{T\to G,i}]\in\mathbb{R}^{2\cdot d_{PLM}}$
@@
(10) $\mathbf{h}_{fused,i}=\text{FusionNet}(\mathbf{h}_{concat,i})\in\mathbb{R}^{d_{fused}}$
@@
(11) $\hat{\mathbf{y}}_{i}=\text{Softmax}(\text{MLP}_{classifier}(\mathbf{h}_{fused,i}))$
@@
5. Training Procedure
The model training proceeds in two distinct stages: (1) GNN Contrastive Pre-training, focused on learning robust structural representations and initializing the GNN-specific mask token, followed by (2) End-to-End Fine-tuning, where the entire model is trained for the downstream node classification task. This two-stage approach allows the GNN to develop a strong structural foundation before being integrated with the powerful but computationally intensive language model. All training was done on a single Nvidia Quadro RTX 8000 (48GB).
5.1. Stage 1: GNN Contrastive Pre-training
@@
5.1.1. Data Augmentation via Soft Masking
@@ 5.1.2. Contrastive Objective
(12) $\mathcal{L}_{pretrain}=\mathcal{L}_{NT-Xent}(\mathbf{H}_{G,1},\mathbf{H}_{G,2},\tau_{pretrain})$
@@
5.2. Stage 2: End-to-End Fine-tuning for Node Classification
@@
5.2.1. Active Node Selection and GNN Perturbation
@@
5.2.2. Supervised Classification Loss
@@
5.2.3. Optimization and Hyperparameters
The total loss $\mathcal{L}_{classify}$ is backpropagated through the entire model. All trainable parameters are updated using the AdamW optimizer (Loshchilov and Hutter, 2019) with differential learning rates:
@@ 5.2.3. Optimization and Hyperparameters
[…]
LLM4HeG (Wu et al., 2025) (fine-tuned LLM/SLMs and distilled SLMs)
[…]
@@
[…]
Table 1. Accuracy for node classification of different methods. Bold denotes the best score while underline denotes the second best.
6. Results
We evaluate GMLM on five widely-used heterophilic benchmark datasets. Following standard protocols, we randomly split the nodes into train, validation, and test sets with a proportion of 48%/32%/20% for the Cornell, Texas, Wisconsin, and Actor datasets, and 50%/25%/25% for Amazon following (Platonov et al., 2024). All reported results in Table 1 are the mean accuracy (%) and standard deviation averaged over 10 independent runs.
@@ 6. Results
A comparison of our model variants reveals that performance scales with the representational power of the PLM, as Snowflake-embed (305M) consistently outperforms GTE-Base (110M). However, even the smaller GMLM-GTE-Base is highly competitive, outperforming the majority of all baselines and securing the second-best result on Wisconsin. This confirms the robustness and effectiveness of the core GMLM framework itself. We provide a full qualitative comparison of the PLM backbones and the effect of different training stages via UMAP visualizations on Cornell, Texas and Wisconsin in Section 7.
7. Effect of Training Stages
7.1. Cornell
We study the effect of different stages of training our GMLM framework with the different PLM backbones. Figure 2 and Figure 3 show the embedding visualizations of the node embeddings on the Cornell dataset, projected into a 2D space using UMAP. The plots illustrate the state of embeddings for various node types (students, courses, etc.) at three key stages: initial raw features, after contrastive pre-training, and after full fine-tuning. Figure 2 presents the results using the GTE-base model as the text encoder. While it shows a noticeable improvement in cluster formation from the initial state to the fine-tuned stage, the final clusters remain somewhat diffuse and exhibit overlap. In contrast, Figure 3, which uses the larger Snowflake embedding model, demonstrates markedly superior performance. Following the full fine-tuning process, the node clusters are significantly more compact, distinct, and well-separated. This visual evidence indicates that the Snowflake model generates more discriminative feature representations and strongly supports the conclusion that it serves as a more powerful text encoder for this particular dataset and task.
Figure 2. UMAP visualization of node embeddings using the GTE-base model (PLM backbone). The three panels show the embeddings for raw features, after contrastive pre-training, and after full fine-tuning.
Figure 3. UMAP visualization of node embeddings using the larger Snowflake embedding model (PLM backbone). These panels show a clear improvement in cluster separation after fine-tuning compared to the GTE-base model in Figure 2.
7.2. Texas
Similarly, we analyze the training effects on the Texas dataset. Figure 4 and Figure 5 display the UMAP visualizations of node embeddings, highlighting the progression from initial features through contrastive pre-training to full fine-tuning. In Figure 4, the GTE-base model, while improving cluster separation compared to the raw features, still results in embeddings where the classes are not sharply delineated after fine-tuning. Conversely, the larger Snowflake embedding model, shown in Figure 5, yields significantly better results. After the full fine-tuning stage, the node clusters are dense and clearly separated, demonstrating the superior capability of the Snowflake model to learn discriminative representations for the node types within the Texas dataset. This again underscores the impact of the PLM backbone’s scale and architecture on the final performance of the GMLM framework.
Figure 4. UMAP visualization of node embeddings on the Texas dataset using the GTE-base model. The plots show the progression from raw features to fine-tuned embeddings.
Figure 5. UMAP visualization of node embeddings on the Texas dataset using the Snowflake embedding model. A clear improvement in cluster quality is visible compared to the GTE-base model shown in Figure 4.
Figure 6. UMAP visualization of node embeddings on the Wisconsin dataset using the GTE-base model. The panels show the progression from raw features to fine-tuned embeddings, resulting in moderately separated clusters.
Figure 7. UMAP visualization of node embeddings on the Wisconsin dataset using the Snowflake embedding model. A dramatic improvement in cluster density and separation is visible compared to the GTE-base model shown in Figure 6.
7.3. Wisconsin
Finally, we present the results for the Wisconsin dataset. The visualization of node embeddings in Figure 6 and Figure 7 follows the same three-stage training progression. For the GTE-base model (Figure 6), the fine-tuning process yields some structure, but the resulting clusters lack clear separation and show considerable overlap between different node types. In stark contrast, the Snowflake embedding model (Figure 7) demonstrates exceptional performance. The clusters corresponding to different node types are highly compact and distinctly separated after full fine-tuning. This comparison on the Wisconsin dataset further reinforces our finding that the larger PLM backbone is crucial for learning high-quality, discriminative node representations within our GMLM framework.
8. Conclusion
@@
In this paper, we introduced the Graph Masked Language Model (GMLM), a novel framework that effectively integrates deep graph structural learning with the rich semantic understanding of Pre-trained Language Models. By employing a two-stage training procedure featuring a contrastive GNN pre-training phase and an end-to-end fine-tuning stage with dynamic active node selection, our model addresses key challenges of scalability and robust representation learning. The use of a sophisticated bi-directional cross-attention module for fusion proved critical, enabling GMLM to achieve new state-of-the-art results on four out of five widely-used heterophilic benchmarks. Our model significantly outperforms both specialized GNNs and much larger LLM-based approaches, underscoring that a carefully designed, deeply integrated architecture can be more powerful and capital-efficient than brute-force scaling for text-rich graph representation learning.
9. Limitations and Future Work